A small firm has an ending inventory of $52,000 as at December 31, 2012 and the following accounting information.

Month

Ending Inventory

Cost of Goods Sold

January 2013

$75,000

$225,000

February

$56,000

$325,000

March

$25,000

$240,000

April

$85,000

$325,000

May

$125,000

$460,000

June

$95,000

$220,000

July

$72,000

$85,000

August

$45,000

$156,000

September

$52,500

$220,000

October

$120,000

$265,000

November

$162,500

$100,000

December

$255,000

$350,000

a) Compute the monthly inventory turnover ratio for each of the twelve months. Do you see any trend in the monthly inventory turnover ratio? (4 points)

Hints: the average inventory level = (ending inventory of previous period + ending inventory of current period)/2. Use Excel to do all calculations and copy/paste the result to Word file.

b) What are the annual cost of goods sold and the average inventory for the year? (3 points)

Hint: the annual cost of goods sold = summation of cost of goods sold of 12 months; the average inventory for the year = summation of average inventory of 12 months / 12.

c) Compute the annual inventory turnover ratio. What can the purchasing department do to improve the firm’s performance? (3 points)

Problem 2 (10 points)

You are given the following information:

Costs

Make Option

Buy Option

Fixed Cost

$25,000

$3,000

Variable Cost

$8

$12

  1. Find the break-even quantity and the total cost at the break-even point. (4 points)
  2. If the requirement is 4,500 units, is it more cost-effective for the firm to buy or make the components? What is the cost savings for choosing the cheaper option? (3 points)
  3. If the requirement is 6,000 units, is it more cost-effective for the firm to buy or make the components? What is the cost savings for choosing the cheaper option? (3 points)
Answer #1

1.
a. Inventory Turnover = Cost of Goods Sold / Average Inventory

Month Ending Inventory Cost of Goods Sold Average Inventory Inventory Turnover
Jan-13 75,000 2,25,000 63,500 3.54
February 56,000 3,25,000 65,500 4.96
March 25,000 2,40,000 40,500 5.93
April 85,000 3,25,000 55,000 5.91
May 1,25,000 4,60,000 1,05,000 4.38
June 95,000 2,20,000 1,10,000 2.00
July 72,000 85,000 83,500 1.02
August 45,000 1,56,000 58,500 2.67
September 52,500 2,20,000 48,750 4.51
October 1,20,000 2,65,000 86,250 3.07
November 1,62,500 1,00,000 1,41,250 0.71
December 2,55,000 3,50,000 2,08,750 1.68
Total 29,71,000 10,66,500


b. Annual Cost of Goods Sold = $2,971,000
Annual Average Inventory = $1,066,500 / 12 = $88,875

c. Annual Inventort Turnover = $2,971,000 / $88,875 = 33.43 times
To improve the same, company should try to increase the sales.

Problem 2
Break Even Quantity = ($25000-3000)/(12-8) = 5500
Cost at Break Even Point = 5500 x $8 + $25000 = $69,000

For 4500 units,
Buy Option should be chosen, as quantity is lower than break even point.
Cost Savings = ($25000+4500*8) - ($3000+4500*12) = $4000

For 6000 units,
Make Option should be chosen, as quantity is greater than break even point.
Cost Savings = ($3000+6000*12) - ($25000+6000*8) = $2000